OpenStudy (anonymous):

Simplify this expression. a9 × a18 = A. a2 B. a9 C. a27 D. a162

OpenStudy (diyadiya):

\[\large x^m \times x^n = x^{m+n}\]

OpenStudy (diyadiya):

\[a^9 \times a^18= a^{9+18}\]

OpenStudy (anonymous):

thanks so it would be C. a27

OpenStudy (diyadiya):

Yes (: